# The neutrino luminosity and energy spectrum of nova outburst  [^1] Released on March, 1st, 2021

### v1

@@ 3 Result
Figure 8: The neutrino luminosity curves of T CrB nova. The left panel is first two bursts of the T CrB nova and the next one shortly after, the right panel is predicted duration of the next outburst of T CrB.
3.4 neutrino energy spectrum
@@
In this section, we consider the neutrino spectrum is produced in two ways by novae, one by nuclear reactions inside novae and the other by hadronic processes outside novae.
3.4.1 nuclear reaction
@@
Considering that the scene of nova outburst is similar to that of supernovae explosion. The CNO cycle that dominates nova outbursts and the fast electron-neutrino bursts $(\mathrm{p}+\mathrm{e}^{-}\longrightarrow\mathrm{n}+v_{\mathrm{e}})$ in supernovae are both beta decay, so we think that the neutrino spectrum produced by the nova nuclear reaction is similar to the supernova neutrino spectrum. Following the Baxter et al. (2022), the initial neutrino energy spectrum has the form:
@@
where $D^{2}_{\mathrm{TE}}$ is the distance between T CrB and Earth, ${\left\langle E_{\nu}\right\rangle}$ is mean energy. The $\alpha$ is spectral shape parameter.
As discussed in section 3.3, the low-energy nuclear neutrino luminosity of the next outburst of T CrB is $2.70\times 10^{8}\ L_{\nu,\odot}$. The CNO cycle produces neutrinos with mean energies of about 2 MeV (Borexino Collaboration et al., 2020). The $\alpha$ is 2.5 (Keil et al., 2003). The electron neutirno spectrum is shown in Figure 9. The model predicts the electron neutrino flux from nuclear reaction of T CrB outburst that can not exceeds the sensitivity limit of Super-Kamiokande.
Figure 9: The estimated total electron neutrino flux reaching the Earth from T CrB (orange line). The Super-Kamiokande’s upper limit on electron neutrino flux (Suzuki, 2019) is blue line.
3.4.2 shock speed
Protons accelerated by the shock of the nova interact with ambient particles in RG wind can produce high-energy neutrinos. The flux of muon neutrinos reaching Earth from the T CrB is calculated using the semi-analytical formulation developed in Kelner et al. (2006).
@@
where c is light speed. $F_{\nu}\left(x,E_{\nu}/x\right)$ represents a neutrino spectrum produced by a single energy proton, $x$ is $E_{\nu}/E_{p}$. $\sigma_{\mathrm{inel}}$ is total cross section of $pp$ interactions
$\sigma_{\rm inel}(E_{\mathrm{p}})=34.3+1.88L+0.25L^{2}\ mb\ ,$ (19)
@@
$\alpha$ is power-law index which we select 2.2, $\beta=1$ (Kelner et al., 2006). A is determined based on the given condition
@@
where the minimum proton energy is the the rest mass energy of proton ($E_{\mathrm{min}}\approx 1\rm\ GeV$). The $E_{\rm{dens,T\ CrB}}$ is the energy density of protons around T CrB, it is determined by the kinetic energy of the shock wave (Acciari et al., 2022):
@@
where $M_{\mathrm{ej}}$ is the total ejection mass during the eruption. Based on our T CrB nova model, $M_{\mathrm{ej}}=1.36\times 10^{-6}\ \rm M_{\odot}$. According to Caprioli and Spitkovsky (2014), the $10\%$ of shock kinetic energy can be converted to the energy of high-energy protons. For T CrB nova, the energy released into accelerated protons ($E_{\rm p,nova}$) is about $2.72\times 10^{43}\ \rm erg$. Meanwhile, the energy density of protons surrounding the T CrB can be determined by dividing the total energy by the volume of the corresponding region:
@@
Finally, we can calculate that the proton energy density near the shock wave is $102\ \rm erg/cm^{3}$, which is much greater than that the average energy density in CRs in the Milky Way ( $\sim\ 1\ \rm erg/cm^{3}$).
To validate the credibility of our hadron model, we conduct an estimation on the gamma rays emitted during the RS Oph nova event and compare our results with those presented in De Sarkar et al. (2023). Protons accelerated by the shock of the nova interacting with ambient particles in RG wind also can produce high-energy gamma-ray, $pp$ interactions resulting in the generation of neutral muons, these muons subsequently decay into high-energy gamma-ray ($\pi^{0}\to 2\gamma$). The fitting formula of gamma-ray flux produced by $pp$ interaction has the same form as equation (19):
@@
here $D_{\rm RE}$ is the distance between RS Oph and Earth, $x$ is $E_{\gamma}/E_{p}$. The key parameters for calculating the gamma-ray emissions of RS Oph include : $v_{\mathrm{sh}}=\rm 4500\ km/s$, $r_{\mathrm{sh}}=\rm 4\times 10^{13}\ cm$, $v_{\mathrm{RG}}=\rm 10\ km/s$, $\dot{M}_{\mathrm{RG}}=5\times 10^{-7}\ \rm M_{\odot}\ \mathrm{yr}^{-1}$, $\tau_{pp}\approx 2.2\times 10^{6}\ \mathrm{s}$, $B=0.11\ \rm G$, $\alpha=2.2$, $E_{0}=400\ \rm GeV$. The gamma-ray flux from RS Oph to Earth is depicted in Figure 10. Our analysis reveals that the MAGIC and H.E.S.S. gamma-ray data points can be consistently explained by gamma-ray generated within our hadronic model.
@@
Figure 10: The estimated gamma-ray flux produced by by the $pp$ interaction from the RS Oph outburst. The orange line is the gamma-ray flux produced by the $pp$ interaction from the RS Oph outburst predicted by De Sarkar et al. (2023). The blue line is the total muonic neutrino flux from the RS Oph outburst by our hadronic model. H.E.S.S. and MAGIC data points are taken from Aharonian et al. (2022), and Acciari et al. (2022), respectively.
@@
The muon neutrino flux after considering neutrino oscillations is illustrated in Figure 11. Neutrino oscillations that its flavor states (electron neutrinos, muon neutrinos, and tau neutrinos) change during propagation. That is, the muon neutrino flux reaching Earth is 1/3 of the total neutrino flux produced by the $pp$ interactions at the T CrB. The blue line is the neutrino flux from the RS Oph outburst predicted by De Sarkar et al. (2023), we reproduce their result and compare it with our predictions of the neutrino flux of the T CrB. The neutrino spectrum generated by the T CrB outburst exhibits similarities to that of RS Oph, as predicted by Gagliardini et al. (2024). However, due to its closer proximity and higher maximum energy of proton production during the outburst, the resulting neutrino flux reaching Earth is amplified. As shown in Figure 11, it is unlikely that the current-generation IceCube will observe it.
@@
Figure 11: The estimated total muonic neutrino flux reaching the Earth from T CrB (orange line). The blue line is the total muonic neutrino flux from the RS Oph outburst. The IceCube’s upper limit on muon neutrino flux is green line (Pizzuto et al., 2021).
4 Conclusions
@@
Using MESA, we simulate the evolution for nova outburst with CO WD mass ranging from 0.6 to 1.1 $\rm M_{\odot}$, and 1.1 $\rm M_{\odot}$ ONe WD. We calculate the low-energy nuclear and thermal neutrino luminosity, and give neutrino H-R diagram as well as low-energy nuclear and thermal neutrino luminosity curves. Our results show: During the accretion phase, the nuclear neutrino is mainly produced by the pp chains, and the thermal neutrino is mainly dominated by plasma decay. Neutrinos are produced mainly by thermal processes. Compared with photons, neutrinos carry away less energy. However, during the TNR, the temperature of the nuclear reaction zone on the surface of the nova exceeds $7\times 10^{7}\rm\ K$, the nuclear neutrino is mainly dominated by the CNO cycle, and the thermal neutrino is mainly dominated by photon neutrino, nuclear neutrino will far more exceed thermal neutrino. Meanwhile, photon is easily trapped during TNR, the low-energy nuclear and thermal neutrino luminosity rises rapidly and is much larger than the photon luminosity.
@@
The $M_{\mathrm{WD}}$, $\dot{M}$, $T_{\mathrm{C}}$, and mixing have significant effects on the low-energy nuclear neutrino luminosity of nova. The temperature of the nuclear reaction zone on the surface of a WD increases with mass, so does the low-energy nuclear neutrino luminosity. The smaller the $\dot{M}$, the longer the outburst interval, the more accreted material, and the more intense the outburst, so the greater the low-energy nuclear neutrino luminosity. The model that does not take into account mixing has a higher low-energy nuclear neutrino luminosity because it has a denser accretion layer and a higher temperature. Cold nova model has lower low-energy neutrino luminosity in the accretion phase and higher low-energy nuclear neutrino luminosity during the TNR. It is because its core and shell are both cooler during the accretion phase. And it takes longer to reach the TNR, so more material is accreted, the shell is hotter, and the eruptions are more intense.
@@
We predict that the low-energy nuclear neutrino peak luminosity of the next outburst of the T CrB will be $2.7\times 10^{8}\ L_{\nu,\odot}$ and a low-energy nuclear neutrino outburst duration of 88 days. We also predict the low-energy nuclear and high-energy hadronic neutrino flux that will reach Earth during the next outburst of T CrB. The results show that it is unlikely for current-generation Icecube to observe the neutrino flux produced from the hadronic processes during the T CrB next outburst.
Acknowledgements
@@
This work received the generous support of the National Natural Science Foundation of China under grants No.12163005, U2031204, 12373038, 12288102 and 12263006, the science research grants from the China Manned Space Project with No. CMSCSST-2021-A10, and the Natural Science Foundation of Xinjiang No. 2022D01D85 and 2022TSYCLJ0006, and the Major Science and Technology Program of Xinjiang Uygur Autonomous Region under grant No.2022A03013-3.
@@
The data generated, analysed and presented in this study are available from the corresponding authors on reasonable request.

### v2

@@ III.3 Neutrino luminosity of T CrB nova
Figure 8: The neutrino luminosity (left) and photon luminosity (right) curves of T CrB nova. Left: the first two bursts of the T CrB nova and the next one shortly after. Right: predicted duration of the next outburst of T CrB.
III.4 Neutrino energy spectrum
@@
In this section, we consider the neutrino spectrum is produced in two ways by novae: one by nuclear reactions inside novae and the other by hadronic processes outside novae.
III.4.1 Nuclear reaction
@@
Consider that the scene of the nova outburst is similar to that of supernovae explosion. The CNO cycle that dominates nova outbursts and the fast electron-neutrino bursts $(\mathrm{p}+\mathrm{e}^{-}\longrightarrow\mathrm{n}+v_{\mathrm{e}})$ in supernovae are both $\beta$ decay, so we think that the neutrino spectrum produced by the nova nuclear reaction is similar to the supernova neutrino spectrum. Following Baxter et al. [90], the initial neutrino energy spectrum has the form
@@
where $D^{2}_{\mathrm{TE}}$ is the distance between T CrB and Earth, and ${\left\langle E_{\nu}\right\rangle}$ is mean energy. The $\alpha$ is a spectral shape parameter.
As discussed in Sec. III.3, the low-energy nuclear neutrino luminosity of the next outburst of T CrB is $2.70\times 10^{8}\ L_{\nu,\odot}$. The CNO cycle produces neutrinos with mean energies of about 2 MeV [91]. The $\alpha$ is 2.5 [92]. The electron neutrino spectrum is shown in Fig. 9. The model predicts the electron neutrino flux from nuclear reaction of T CrB outburst that cannot exceed the sensitivity limit of Super-Kamiokande.
Figure 9: The estimated total electron neutrino flux reaching Earth from T CrB (orange line). The Super-Kamiokande’s upper limit on electron neutrino flux [93] is the blue line.
III.4.2 Shock speed
Protons accelerated by the shock of the nova interact with ambient particles in the RG wind can produce high-energy neutrinos. The flux of muon neutrinos reaching Earth from the T CrB is calculated using the semianalytical formulation developed in Kelner et al. [94],
@@ III.4.2 Shock speed
where c is light speed. $F_{\nu}\left(x,E_{\nu}/x\right)$ represents a neutrino spectrum produced by a single energy proton, and $x$ is $E_{\nu}/E_{p}$. $\sigma_{\mathrm{inel}}$ is total cross section of $pp$ interactions,
$\sigma_{\rm inel}(E_{\mathrm{p}})=34.3+1.88L+0.25L^{2}\rm\ mb\ ,$ (19)
@@
where $\alpha$ is the power-law index, which we select to be 2.2, and $\beta=1$ [94]. $A$ is determined based on the given condition:
@@
where the minimum proton energy is the rest mass energy of the proton ($E_{\mathrm{min}}\approx 1\rm\ GeV$). The $E_{\rm{dens,T\ CrB}}$ is the energy density of protons around T CrB, determined by the kinetic energy of the shock wave [39],
@@
where $M_{\mathrm{ej}}$ is the total ejection mass during the eruption. Based on our T CrB nova model, $M_{\mathrm{ej}}=1.36\times 10^{-6}\ \rm M_{\odot}$. According to Caprioli and Spitkovsky [95], $10\%$ of shock kinetic energy can be converted to the energy of high-energy protons. For the T CrB nova, the energy released into accelerated protons ($E_{\rm p,nova}$) is about $2.72\times 10^{43}\ \rm erg$. Meanwhile, the energy density of protons surrounding the T CrB can be determined by dividing the total energy by the volume of the corresponding region,
@@
Finally, we can calculate that the proton energy density near the shock wave is $102\ \rm erg/cm^{3}$, which is much greater than that of the average energy density in cosmic rays in the Milky Way ( $\sim\ 1\ \rm erg/cm^{3}$).
To validate the credibility of our hadron model, we conduct an estimation on the $\gamma$ rays emitted during the RS Oph nova event and compare our results with those presented by De Sarkar et al. [40]. Protons accelerated by the shock of the nova interacting with ambient particles in RG wind also can produce high-energy $\gamma$-ray, $pp$ interactions resulting in the generation of neutral muons; these muons subsequently decay into high-energy $\gamma$-ray ($\pi^{0}\to 2\gamma$). The fitting formula of $\gamma$-ray flux produced by $pp$ interaction has the same form as Eq. (18),
@@
where $D_{\rm RE}$ is the distance between RS Oph and Earth, and $x$ is $E_{\gamma}/E_{p}$. The key parameters for calculating the $\gamma$-ray emissions of RS Oph include $v_{\mathrm{sh}}=\rm 4500\ km/s$, $r_{\mathrm{sh}}=\rm 4\times 10^{13}\ cm$, $v_{\mathrm{RG}}=\rm 10\ km/s$, $\dot{M}_{\mathrm{RG}}=5\times 10^{-7}\ \rm M_{\odot}\ \mathrm{yr}^{-1}$, $\tau_{pp}\approx 2.2\times 10^{6}\ \mathrm{s}$, $B=0.11\ \rm G$, $\alpha=2.2$, and $E_{0}=400\ \rm GeV$. The $\gamma$-ray flux from RS Oph to Earth is depicted in Fig. 10. Our analysis reveals that the MAGIC and H.E.S.S. $\gamma$-ray data points can be consistently explained by $\gamma$ rays generated within our hadronic model.
@@
Figure 10: The estimated $\gamma$-ray flux produced by the $pp$ interaction from the RS Oph outburst. The orange line is the $\gamma$-ray flux produced by the $pp$ interaction from the RS Oph outburst predicted by De Sarkar et al. [40]. The blue line is the total muonic neutrino flux from the RS Oph outburst by our hadronic model. H.E.S.S. and MAGIC data points are taken from Aharonian et al. [78] and Acciari et al. [39], respectively.
@@
The muon neutrino flux after considering neutrino oscillations is illustrated in Fig. 11. Neutrino oscillations refer to the phenomenon where neutrinos change from one flavor state to another during propagation, including transformations between electron neutrinos, muon neutrinos, and $\tau$ neutrinos. That is, the muon neutrino flux reaching Earth is 1/3 of the total neutrino flux produced by the $pp$ interactions at the T CrB. The blue line is the neutrino flux from the RS Oph outburst predicted by De Sarkar et al. [40]. We reproduce their result and compare it with our predictions of the neutrino flux of the T CrB. The neutrino spectrum generated by the T CrB outburst exhibits similarities to that of RS Oph, as predicted by Gagliardini et al. [96]. However, due to its closer proximity and higher maximum energy of proton production during the outburst, the resulting neutrino flux reaching Earth is amplified. As shown in Fig. 11, it is unlikely that the current-generation IceCube will observe it.
@@
Figure 11: The estimated total muonic neutrino flux reaching Earth from T CrB (orange line). The blue line is the total muonic neutrino flux from the RS Oph outburst. IceCube’s upper limit on muon neutrino flux is the green line [97].
IV Conclusions
@@
Using mesa, we simulate the evolution for nova outburst with CO WD mass ranging from 0.6 to 1.1 $\rm M_{\odot}$, and 1.1 $\rm M_{\odot}$ ONe WD. We calculate the low-energy nuclear and thermal neutrino luminosity and give neutrino H-R diagram as well as low-energy nuclear and thermal neutrino luminosity curves. Our results show that, during the accretion phase, the nuclear neutrino is mainly produced by the pp chains, and the thermal neutrino is mainly dominated by plasma decay. Neutrinos are produced mainly by thermal processes. Compared with photons, neutrinos carry away less energy. However, during the TNR, the temperature of the nuclear reaction zone on the surface of the nova exceeds $7\times 10^{7}\rm\ K$, the nuclear neutrino is mainly dominated by the CNO cycle, and the thermal neutrino is mainly dominated by photon neutrino, and the nuclear neutrino will far exceed the thermal neutrino. Meanwhile, the photon is easily trapped during TNR, and the low-energy nuclear and thermal neutrino luminosity rises rapidly and is much larger than the photon luminosity.
@@
The $M_{\mathrm{WD}}$, $\dot{M}$, $T_{\mathrm{C}}$, and mixing have significant effects on the low-energy nuclear neutrino luminosity of novae. The temperature of the nuclear reaction zone on the surface of a WD increases with mass, so does the low-energy nuclear neutrino luminosity. The smaller the $\dot{M}$, the longer the outburst interval, the more accreted material, and the more intense the outburst, so the greater the low-energy nuclear neutrino luminosity. The model that does not take into account mixing has a higher low-energy nuclear neutrino luminosity because it has a denser accretion layer and a higher temperature. The cold nova model has lower low-energy neutrino luminosity in the accretion phase and higher low-energy nuclear neutrino luminosity during the TNR. This is because its core and shell are both cooler during the accretion phase. It takes longer to reach the TNR, so more material is accreted, the shell is hotter, and the eruptions are more intense.
@@
We predict that the low-energy nuclear neutrino peak luminosity of the next outburst of the T CrB will be $2.7\times 10^{8}\ L_{\nu,\odot}$ with a low-energy nuclear neutrino outburst duration of 88 days. We also predict the low-energy nuclear and high-energy hadronic neutrino flux that will reach Earth during the next outburst of T CrB. The results show that it is unlikely for current-generation IceCube to observe the neutrino flux produced from the hadronic processes during the T CrB next outburst.
Acknowledgements.
@@
This work received the generous support of the National Natural Science Foundation of China under Grants No. 12163005, No. U2031204, No. 12373038, No. 12288102, and No. 12263006, the science research grants from the China Manned Space Project with No. CMSCSST-2021-A10, the Natural Science Foundation of Xinjiang No. 2022D01D85 and No. 2022TSYCLJ0006, and the Major Science and Technology Program of Xinjiang Uygur Autonomous Region under Grant No. 2022A03013-3.
@@
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.